OpenStudy (anonymous):

How do I solve this series using the first 5 terms? SERIES BELOW!

OpenStudy (anonymous):

Its the ramanujan series for estimating pi \[\frac{ 2\sqrt{2} }{ 9801} \sum_{k=0}^{\infty} \frac{ (4k!)(1103+26390k) }{ k!^{4}* 396^{4k} }\]

OpenStudy (anonymous):

Don't get to particular cuz i'm technically writing a computer program for this. I just don't really understand the math it is asking for.

OpenStudy (anonymous):

If k = 5, do i plug in the values 0-5 in the right side of the equation adding each answer up as I go and then multiply by the fraction out front? Or do I multiply by the fraction out front each time I plug k in?

OpenStudy (anonymous):

figured it out. thanks!